A bus moving at a speed of 20m/s begins to slow at a constant rate of 3.0m/s each second. How far does it go before stopping?


1
Expert's answer
2021-03-29T16:47:35-0400

Let's first find the time that the bus takes to stop:


"t=\dfrac{v-v_0}{a}=\dfrac{0-20\ \dfrac{m}{s}}{-3.0\ \dfrac{m}{s^2}}=6.67\ s."

Finally, we can find the distance travelled by the bus before it stops:


"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\cdot(20\ \dfrac{m}{s}+0)\cdot6.67\ s=66.7\ m."